# Understanding Radiation Flow in a Stochastic Medium

### v1

@@ Abstract
Radiation flow through an inhomogeneous medium is critical in a wide range of physics and astronomy applications from transport across cloud layers on the earth to the propagation of supernova blast-waves producing UV and X-ray emission in supernovae. This paper reviews the current state of the art in the modeling of inhomogeneous radiation transport, subgrid models developed to capture this often-unresolved physics, and the experiments designed to improve our understanding of these models. We present a series of detailed simulations (both single-clump and multi-clump conditions) probing the dependence on the physical properties of the radiation front (e.g. radiation energy) and material characteristics (specific heat, opacity, clump densities). Unless the radiation pressure is high, the clumps will heat and then expand, effectively cutting off the radiation flow. The expanding winds can also produce shocks that generates high energy emission. We compare our detailed simulations with some of the current subgrid prescriptions, identifying some of the limitations of these current models.
@@ 1 Introduction
For many physics and astrophysics applications, energy transport is often carried by photons, neutrinos and electrons. In its generality, solving this energy transport problem requires modeling a seven dimensional problem where the spatial, angular and energy distribution of these carriers must be evolved (see reviews https://www.osti.gov/servlets/purl/917504). Coupling this energy transport to matter drives further hydrodynamic instabilities, pushing the limits of our current modeling capabilities. This has led to many approximate methods in modeling radiation-hydrodynamics that include simplifications in the radiation transport (e.g. integrating over angle or energy), coupling terms (e.g. modeling pure transport or pure hydrodynamics models), or spatial dimensions (homogenizing the radiation flow into a 1-dimensional flow). These simplifying assumptions are all valid in specific conditions. For example, for supernova shocks traveling within a star, radiation is trapped in the flow, but after the shock breaks out of the star well into the circumstellar medium, radiation flows freely from where it is emitted and the hydrodynamics can be approximated by a simple homologous outflow. In stars, although the radiation is not trapped, it is in the “diffusive” regime where the angular distribution is well-defined and integrating over angle can remove dimensions from the numerical problem.
Although many studies exist that investigate different approximations in radiation transport or radiation-hydrodynamics coupling, there are far fewer studies of the role of inhomogeneities in energy transport (also known as heterogeneous transport or transport through a clumpy medium). The effect of a clumpy medium can modify transport solutions across a broad range of approximation regimes. In transport-only solutions, inhomogeneous media can alter the angular distribution of the radiation. In optically thin applications, these more optically-thick clumps screen the photons, reducing the radiative flux. In more optically-thick regimes, they alter the radiation flow. In radiation-hydrodynamics calculations, a broad range of additional physical effects from radiation/clump interactions can alter the radiative flow. Radiation can cause ablation or, through heating, drive a wind from the clumps. In this paper, we refer to ablation when the disruption of the material is due to momentum deposition and “blow-off” or “wind” when the outflow is due to energy injection. This can inject higher density and, in some cases, higher opacity material into the surroundings, increasing the impact of the clumps on the radiation flow. Shocks from these winds can produce high-energy emission, altering the energy spectrum of the radiation. In some cases, radiative pressure can cause the clumps to implode. The importance of these different effects depends on the applications and their conditions.
@@
This paper investigates the radiation flow through these clumpy media, studying the physics effects on the radiation transport and coupling between radiation and hydrodynamics. Understanding the details of radiation flow in clumpy media will enable a better use of observations in nature and results from high energy-density experiments to constrain and validate simulations (codes, numerical methods, physical data, models, and the approximations therein). Although becoming less rare, modeling a clumpy medium at scale severely strains computational resources, even on massive supercomputers. In addition to the aforementioned numerical approximations, simplifications include simulating a small subset of the medium and homogenizing or under-resolving small geometric details. An under-resolved simulation tests the asymptotic behavior of the numerical methods and the accuracy of averaging methods for geometry, materials, equations of state, and opacities. Furthermore, a clumpy medium that we simulate or fabricate for an experiment is usually a single and particular realization of a stochastic medium, which is defined probabilistically. Transport-in-Stochastic-Media methods have been developed to produce ensemble-average linear transport solutions, as if an infinite number of simulations of media sampled from the stochastic definitions were run and averaged. The applications of these methods to coupled radiation-hydrodynamics are much less mature, as is an accompanying estimation of the standard deviation of the ensemble average. For the bulk of this paper, we focus on energy transport via photon radiation. Many of the conclusions also apply to other energy carriers including neutrinos and electrons scaled by the transport regime (mean free path) and emission/deposition properties. Section 2 reviews a set of applications where radiative transport through an inhomogeneous medium is important, discussing which physical effects are most critical in each application. Section 3 reviews the current state of both modeling and experiment to study inhomogeneous radiation flow. In this paper, we focus on detailed radiation-hydrodynamics calculations to better understand the physics behind radiation flow and Section 4.1 describes our initial conditions and the computational methods used in this paper. Section 4.2 focuses on the basic physics behind radiation flow across a single clump. Section 4.3 studies the more complex interactions in a multi-clump medium. We conclude with a discussion of how the physics studied in these simulations affects the different applications.
@@ 2 Applications
Radiation transport is critical in a wide range of applications including oncological treatments, radiation shielding, nuclear reactors, astrophysics, and planetary atmospheres; for a review, see (Byvank et al., 2024).
In some cases (particularly in astrophysics), the deposition of energy and momentum from this radiation plays a critical role in the evolution of the phenomena. For example, momentum deposition from radiation is key in understanding mass outflow from stars. Astronomers are now realizing that the smooth, line-driven winds (Puls et al., 2008) are too simple to explain all of the features in the stellar wind profile. Explosive shell burning, opacity-driven instabilities and pressure waves can all produce bursts of mass ejection in a star, producing inhomogeneities in the circumstellar medium including both shells and clumpy media (Fryer et al., 2006; Herwig et al., 2014; Quataert et al., 2016). Studies of line-driven winds indicate that even these, relatively quiescent, outflows, can produce large inhomogeneities in the circumstellar medium (Owocki & Rybicki, 1984; Puls et al., 2008; Sundqvist et al., 2018; Jiang et al., 2018; Owocki et al., 2019). In particular, clumping in the line-driven wind can alter the radiation and mass outflow from these winds (Sundqvist et al., 2018; Owocki & Sundqvist, 2018). Understanding the interaction between radiation and these clumps determines both this mass loss and the evolution of these clumps.
The clumpy structure produced in stellar winds provides the initial conditions through which supernova explosions propagate. Radiation is typically trapped in the material shock until the supernova breaks out of the star. It is believed that the edge of the star is marked by a sharp density break as the stellar profile transitions to a stellar wind profile. In simple spherically symmetric models, the radiation in the shock rapidly transitions from being trapped in the material flow to freely streaming out of the shock (Waxman & Katz, 2017). This effect depends sensitively on the structure of the edge of the star as well as asymmetries in the supernova blastwave (from the explosion and the stellar envelope) which complicate this picture (Bayless et al., 2015; Lovegrove et al., 2017; Fryer et al., 2020a; Irwin et al., 2021). Especially in Wolf-Rayet stars where shock breakout actually occurs in the wind, not the edge of the star, shock breakout can be drastically modified by the radiation flow through the clumpy wind material (Fryer et al., 2020a). Interactions between the radiation and the clumps disrupt the clumps and the ablation of these clumps produce shocks that drive high-energy emission (Fryer et al., 2020a).
@@
Climate and environmental studies have also studied inhomogeneous radiation flow. Understanding the transport through vegetation canopies is important both for agriculture and climate modeling (Kassianov, 2003; Shabanov et al., 2007). Climate scientists have studied radiative transfer through inhomogeneous cloud cover (Cook et al., 2004; Hogan et al., 2016) and recipes to mimic that transfer are used in many climate codes (Wu, 2013; Hogan et al., 2016). With improving satellite data, tests of these models become increasingly stringent, driving advances in this transport.
@@ 3.1 Inhomogeneous Transport Solutions
The effects of radiation transport through inhomogeneous media depends on the properties of the inhomogeneities. Studies of inhomogeneities generally consider a 2-component medium consisting of a radiation flow region with embedded clumps. Throughout this paper, we will refer to the “flow” region as the lower-density region through which the radiation front propagates, and the clump region as the denser clumps around which the radiation flows. In experiments, the flow region is often composed of a silicate foam and the clumps are inclusions in that foam. In many astrophysical applications, the flow region is simply the lower density region where the clumps are produced by dense turbulent eddies. The effect of the clumps depends on their densities, opacities, specific heats, geometry, sizes, and covering fractions. It depends on the transport regime: diffusion limit, transport regime, free-streaming limit of the radiation flow. In the free-streaming limit, dense clumps can completely alter the nature of the flow. For instance, inhomogeneities in atmospheric clouds can lead to an increase in the albedo of the cloud layer, decreasing the amount of radiation that reaches a planet’s surface. In the diffusive regime, radiation is more able to flow around clumps. Transport-regime conditions lie in between these two extremes.
@@
A straightforward approach to introduce inhomogeneites in reduced order techniques is to derive an effective opacity in the region with inhomogeneities. For example, the approach of Prinja & Olson (2005); Olson (2007) included an additional effective scattering opacity as well as a modification to the absorption opacity for an inhomogeneous medium for a purely absorptive medium. The effective absorption opacity ($\sigma_{\rm a,eff}$) is given by:
@@
$\sigma_{\rm s,eff}=\nu^{2}/[\tilde{\sigma}(1+\lambda_{c}\tilde{\sigma})]$ (2)
where $\tilde{\sigma}=p_{1}\sigma_{a1}+p_{2}\sigma_{a2}$ is the average opacity, $\nu^{2}=p_{1}p_{2}(\sigma_{1}-\sigma_{2})^{2}$, and $\lambda_{c}=\lambda_{1}\lambda_{2}/(\lambda_{1}+\lambda_{2})$ is the geometric mean of the chord lengths, where the chord length is defined by the length of a chord covered by each material.
Using this prescription, we can calculate the effective absorption and scattering opacities as a function of the covering fraction of the clumps. Figure 1 shows the effective absorption and scattering opacities under this formalism as a function of the clump covering fraction. Although the effective absorption cross section does not increase until the clump covering fraction is large, the effective scattering term quickly approaches the value of the clump opacity. This scattering term can drastically lower the propagation speed of the radiation.
Figure 1: Effective absorption (solid) and scattering (dashed) opacities for three values of clump opacity (in ${\rm cm}^{2}\,{\rm g}^{-1}$) as a function of clump covering fraction. The opacity of the flow region is $10{\rm\,cm^{2}\,g^{-1}}$ and the density of the flow, clump regions are $0.05{\rm\,g\,cm^{-3}},5{\rm\,g\,cm^{-3}}$ respectively. The effective absorption opacity does not increase significantly until the covering fraction exceeds 0.8. However, the effective scattering opacity can be high if the clump opacity is high compared to the opacity in the flow region.
For stellar winds, Sundqvist et al. (2014) used a different approach, setting the opacity to
$\sigma=\tilde{\sigma}(1+\tau_{\rm cl}f_{\rm ic})/(1+\tau_{\rm cl})$ (3)
where $\tau_{\rm cl}$ is the clump optical depth and $f_{\rm ic}$ is the density of the inter-clump medium divided by the average density of the medium.
Most of these past studies focused on transport-only solutions or solutions where radiation/hydrodynamic effects can be minimized. As such, the solutions typically only consider the relative opacities of the different clumps. More detailed studies included the effects of the size scale of the clumps (with respect to the photon mean free path). This effect becomes increasingly important as we move from the diffusive to the free-streaming transport regimes.
3.2 Experiments

### v3

# Radiation-Hydrodynamics Effects in an Inhomogeneous Medium

@@ Abstract
Radiation flow through an inhomogeneous medium is critical in a wide range of physics and astronomy applications from transport across cloud layers on the earth to the propagation of supernova blast-waves producing UV and X-ray emission in supernovae. Radiation interacts with matter driving hydrodynamic feedback that further alters the radiation characteristics (energy and angular distribution). This paper reviews the current state of the art in the modeling of inhomogeneous radiation transport, subgrid models developed to capture this often-unresolved physics, and the experiments designed to improve our understanding of these models. This paper focuses on simulations based on upcoming experiments designed to test this physics. We present a series of detailed simulations (both single-clump and multi-clump conditions) probing the dependence on the physical properties of the radiation front (e.g. radiation energy) and material characteristics (specific heat, opacity, clump densities). We find that, unless the radiation pressure is high, the clumps will heat and then expand, effectively cutting off the radiation flow. The expanding winds can also produce shocks that generates high energy emission. We compare our detailed simulations with some of the current subgrid prescriptions, identifying some of the limitations of these current models.
@@ 1 Introduction
For many physics and astrophysics applications, energy transport is often carried by photons, neutrinos and electrons. In its generality, solving this energy transport problem requires modeling a seven dimensional problem where the spatial, angular and energy distribution of these carriers must be evolved. Coupling this energy transport to matter drives further hydrodynamic instabilities, pushing the limits of our current modeling capabilities (Castor, 2004). The difficulties in modeling radiation-hydrodynamics have led to many approximate methods that include simplifications in the radiation transport. Approximations in the angular distribution include moment approaches (e.g. diffusion, variable Eddington factor and tensor methods). The differences between these approaches lie in the order of the moments used. More sophisticated, but typically more computationally expensive, approaches discretize the angle distribution using either spherical harmonics (e.g., $P_{N}$, Reed (1972); Brunner (2000)) or discrete ordinate (e.g., $S_{N}$, Pomraning (1973)) approaches. Finally, Monte-Carlo approaches where particles/photons are represented by a set of packets have the ability to easily implement complex physics (e.g. aspherical scattering kernels, relativistic effects) but tend to be expensive and are hampered by noise from low packet numbers. Because they have the potential to capture the full angular distribution, Monte Carlo and discretized methods are often termed “higher-order” methods.
In astrophysics, most of the radiation-hydrodynamics calculations have focused on moment solutions, the simplest being the diffusion approximation (Castor, 2004). Flux-limiters have been included to approximate the transition region between diffusive and free-streaming regimes (Levermore & Pomraning, 1981; Bowers & Wilson, 1982). Radiation-hydrodynamics using flux-limited diffusion have been widely used in astrophysics, coupled to a broad range of hydrodynamics schemes from smooth particle hydrodynamics (Herant et al., 1994), fixed grid and adaptive mesh refinement codes (Turner & Stone, 2001). Higher-order transport techniques coupled to hydrodynamics are gradually becoming more common. For neutrino transport, astronomers are beginning to implement higher-order methods, developing spherical harmonic (Radice et al., 2013), discrete ordinate (Yamada et al., 1999; Liebendörfer et al., 2001) and Monte-Carlo (Foucart et al., 2018; Miller et al., 2019) approaches (for a review, see Akaho et al., 2021). Many of these methods couple to the hydrodynamics using operator-splitting techniques. Although the supernova light-curve community uses a variety of higher-order transport techniques (Blondin et al., 2022), only a few efforts have used higher-order radiation transport methods coupled to matter (e.g. Fryer et al., 2020a).
In many astrophysical applications, the implementation of the atomic physics can be extremely challenging, especially in systems where lines dominate the opacity. Capturing line opacities, particularly in problems where there is considerable Doppler broadening can be challenging and a number of approaches ranging from pure Sobolev solutions(Sobolev, 1963) to expansion opacities have been developed (Karp et al., 1977; Friend & Castor, 1983; Eastman & Pinto, 1993; Blinnikov, 1996; Wehrse et al., 2003). Line-dominated opacities are extremely important for stellar winds, and this community has developed specialized opacity implementations and transport techniques to capture this physics (Carlberg, 1980; Kee et al., 2016; Sundqvist et al., 2018; Driessen et al., 2019; Lagae et al., 2021) although more-recent approaches have begun to introduce moment transport techniques such as flux-limited diffusion (Moens et al., 2022; Narechania et al., 2025) and discretized angle solutions (Goldberg et al., 2022a, b).
These simplifying assumptions are all valid in specific conditions. For example, in calculating supernova light-curves, the level of fidelity depends on how much detail we would like to learn from the observations. The simplest assumption is that the radiation is trapped when the supernova blastwave is propagating through the star (pure hydrodynamics calculations are sufficient) and, after shock breakout, shocks are unimportant and pure transport calculations are sufficient. For many supernova light-curve properties, these assumptions are good enough. However, the most of the focus in the supernova community has been on the implementation of the opacities including out-of-equilibrium effects (see code comparison in Blondin et al., 2022). But it is becoming increasingly clear that shock interactions play a major role in supernova light-curves (Frey et al., 2013a; De La Rosa et al., 2017; Smith, 2017; Chatzopoulos & Weide, 2019). For the most part, these shock interactions are included using either recipes (e.g. Dessart et al., 2010; De La Rosa et al., 2017; Niblett et al., 2025) or through radiation-hydrodynamics using the first moment (flux-limited diffusion) methods (Blinnikov & Bartunov, 1993; Frey et al., 2013a). But a growing number of codes are adding improved transport schemes to hydrodynamic solutions, e.g. the variable Eddington Tensor code Hydra (Höflich, 2009).
The promise of high-fidelity observations from upcoming ultraviolet missions, UltraSat (Shvartzvald et al., 2024), UVEX (Kulkarni et al., 2021), and potential X-ray missions (Arcodia et al., 2024) mandates increasingly accurate calculations of shock interactions in supernova light-curves. One of the primary effects seen in supernovae is the interaction with the radiation from the supernova with the clumpy circumstellar medium produced in stellar winds, common envelope mass ejection, and explosive stellar burning (for reviews, see Smith, 2017; Fryer et al., 2020a; Niblett et al., 2025). These interactions alter the observed supernovae in a number of ways. For example, shocks at the boundary with clumps heats the ejecta, altering the temperature of the supernova. the radiative heating can cause the denser clumps or shells to expand and this “blow-off” that drives additional shock heating. The higher temperatures achieved because of this shock heating is believed to produce the extended ultraviolet emission in supernovae (Smith, 2017; De La Rosa et al., 2017; Fryer et al., 2020a; Niblett et al., 2025; Bayless et al., 2022). Understanding the role of this shock heating allows scientists to disentangle the heating from 56Ni decay from other sources (shock heating, long-lived central engines).
[…]
This paper investigates the radiation flow through these clumpy media, studying the physics effects on the radiation transport and coupling between radiation and hydrodynamics. These studies will focus on leveraging laboratory experiments to gain insight into the physics. These experiments can provide insight into the nature of blow-off and its effect on the opacity in a radiation flow. Although this is an important part of the physics in astrophysical applications such as supernovae, it does not study all of the needed physics. For example, while velocity broadening and line opacities in a velocity gradient are important in supernovae, pressure broadening is far more important in these experiments. The implementation of the opacities is more straightforward in the experiments. In our study, we focus on the coupling between the radiation and the hydrodynamics and the effects of the blow-off.
In addition to the aforementioned numerical approximations in the transport, radiation-hydrodynamics solutions are typically under-resolved. Capturing the effects of heating and blow-off in a solution requires resolving the clumps. Numerical diffusion in hydrodynamics can both artificially transport material and introduce a viscosity to problems in fluid flow (Castor, 2004). Improved discretization techniques (e.g. techniques that include subgrid information) and higher resolution can minimize the effects of this false diffusion. Transport can also experience an effective numerical diffusion, a.k.a. teleportation, where radiation artificially diffuse through a material. As with hydrodynamic diffusion, these effects can be mitigated with subgrid models, e.g.“tilt” functions (Poëtte & Valentin, 2020).
These resolution effects become more difficult when a radiation flow passes through an inhomogeneous medium where the inhomogeneities or clumps can not be fully resolved. A number of subgrid models have been developed to try to capture the physics of radiation flow through an inhomogeneous medium. But, as we shall see in this paper, these recipes capture only one aspect (the increase in opacity) of the inhomogeneous radiation flow under the assumption that no hydrodynamic feedback occurs. This paper studies the physical effects in of radiation flow through an inhomogeneous medium including the increase in opacity, but also on effects that alter the angular and spectral features of the radiation such as shock heating as the blow-off from the clumps interact with their surroundings. For supernova observations, this blow-off and the shocks may be the most important effect of inhomogeneous radiation flow. We also study the stochasticity of the interactions. These studies are designed to indentify the that must be captured in a subgrid model to fully incorporate the effects of radiation flow in an inhomogeneous medium.
@@
Although we focus on photon radiation-transport in this paper, manny of the conclusions also apply to other energy carriers including neutrinos and electrons scaled by the transport regime (mean free path) and emission/deposition properties. Section 2 reviews a set of applications where radiative transport through an inhomogeneous medium is important, discussing which physical effects are most critical in each application. Section 3 reviews the current state of both modeling and experiment to study inhomogeneous radiation flow. In this paper, we focus on detailed radiation-hydrodynamics calculations to better understand the physics behind radiation flow and Section 4.1 describes our initial conditions and the computational methods used in this paper. Section 4.2 focuses on the basic physics behind radiation flow across a single clump. Section 4.3 studies the more complex interactions in a multi-clump medium. We conclude with a discussion of how the physics studied in these simulations affects the different applications.
@@ 2 Applications
Radiation transport is critical in a wide range of applications including oncological treatments, radiation shielding, nuclear reactors, astrophysics, and planetary atmospheres; for a review, see (Byvank et al., 2024). In some cases (particularly in astrophysics), the deposition of energy and momentum from this radiation plays a critical role in the evolution of the phenomena. For example, momentum deposition from radiation is key in understanding radiation-driven outflow from massive stars. Astronomers are now realizing that the smooth, line-driven winds (Puls et al., 2008) are too simple to explain all of the features in the stellar wind profile. Explosive shell burning, opacity-driven instabilities and pressure waves can all produce bursts of mass ejection in a star, producing inhomogeneities in the circumstellar medium including both shells and clumpy media (Fryer et al., 2006; Herwig et al., 2014; Quataert et al., 2016). Studies of line-driven winds indicate that even these, relatively quiescent, outflows, can produce large inhomogeneities in the circumstellar medium (Owocki & Rybicki, 1984; Puls et al., 2008; Sundqvist et al., 2018; Jiang et al., 2018; Owocki et al., 2019). In particular, clumping in the line-driven wind can alter the radiation and mass outflow from these winds (Sundqvist et al., 2018; Owocki & Sundqvist, 2018). Understanding the interaction between radiation and these clumps determines both this mass loss and the evolution of these clumps. Especially in line-driven winds, the implementation of the opacity, can play a critical role in determining the effects of these flows. This paper will focus on transport effects and not on opacity implementations. Castor (2004) reviews these methods in detail.
The clumpy structure produced in stellar mass ejection provides the initial conditions through which supernova explosions propagate. Radiation is typically trapped in the material shock until the supernova breaks out of the star. It is believed that the edge of the star is marked by a sharp density break as the stellar profile transitions to a stellar wind profile. In simple spherically symmetric models, the radiation in the shock rapidly transitions from being trapped in the material flow to freely streaming out of the shock (Waxman & Katz, 2017). This effect depends sensitively on the structure of the edge of the star as well as asymmetries in the supernova blastwave (from the explosion and the stellar envelope) which complicate this picture (Bayless et al., 2015; Lovegrove et al., 2017; Fryer et al., 2020a; Irwin et al., 2021). Especially in Wolf-Rayet stars where shock breakout actually occurs in the circumstellar medium, not the edge of the star, shock breakout can be drastically modified by the radiation flow through the clumpy wind material (Fryer et al., 2020a). Interactions between the radiation and the clumps disrupt the clumps and the ablation of these clumps produce shocks that drive high-energy emission (Fryer et al., 2020a).
@@
Climate and environmental studies have also studied inhomogeneous radiation flow. Understanding the transport through vegetation canopies is important both for agriculture and climate modeling (Kassianov, 2003; Shabanov et al., 2007; Ganapol et al., 1999). Climate scientists have studied radiative transfer through inhomogeneous cloud cover (Cook et al., 2004; Hogan et al., 2016) and recipes to mimic that transfer are used in many climate codes (Wu, 2013; Hogan et al., 2016). With improving satellite data, tests of these models become increasingly stringent, driving advances in this transport.
@@ 3.1 Inhomogeneous Transport Solutions
The effects of radiation transport through inhomogeneous media depends on the properties of the inhomogeneities. Studies of inhomogeneities generally consider a 2-component medium consisting of an optically-thin radiation flow region with optically-thick embedded clumps.  [^1] Although not strictly appropriate for radition-driven winds (Moens et al., 2022; Debnath et al., 2024), the embedded clumps do work for models where clumpy hydrogen material from a common envelope mass-ejection phase is surrounded by a subsequent radiatively-driven wind phase. Developing a subgrid model to capture the physics of this radiation flow depends upon the exact physics that must be captured for a given application. Here we discuss three different physical effects:
  • Radiation Propagation: The clumps will alter the propagation of the radiation front. If the optically-thin region is in the free-streaming regime, the optically-thick clumps will reduce the radiation that propagates through this region. If the optically-thin region is close to the transport regime, the clumps can cause the radiation front to decelerate.
  • Radiation Direction: The clumps not only affect the propagation, but also the angular distribution of the radiation as the clumps heat and re-emit. This effect is important for applications but only noticed in codes that capture this physics, e.g. variable Eddington tensor, spherical harmonics ($P_{N}$), discrete ordinate ($S_{N}$), or Monte Carlo schemes.
  • Hydrodynamic Feedback and Shocks: When the clumps are heated, they will expand. In supernovae, the shocks produced by these expanding shocks are critical in understanding early-time ultraviolet and X-ray emission including shock breakout as well as the peak optical emission (Garnavich et al., 2016; Smith, 2017; De La Rosa et al., 2017; Fryer et al., 2020a).
Ideally, any subgrid model would capture all of this physics. But, as we shall see, most models have focused on radiation propagation and, to a lesser extent, angular distribution of the radiation. By not including the hydrodynamic feedback, these methods can underestimate the opacity, preventing an accurate model for the radiation propagation (see Section 4.3). To our knowledge, no subgrid model includes the radiation-induced shocks and the subsequent shock heating.
Here we review a couple of the current subgrid models. Throughout this paper, we will refer to the “flow” region as the optically-thin region through which the radiation front propagates, and the clump region as the denser, optically-thick clumps around which the radiation flows. In experiments, the flow region is often composed of a silicate foam and the clumps are inclusions in that foam. In many astrophysical applications, the flow region is simply the lower density region where the clumps are produced by dense turbulent eddies. The effect of the clumps depends on their densities, opacities, specific heats, geometry, sizes, and covering fractions. It depends on the transport regime: diffusion limit, transport regime, free-streaming limit of the radiation flow. In the free-streaming limit, dense clumps can completely alter the nature of the flow. For instance, inhomogeneities in atmospheric clouds can lead to an increase in the albedo of the cloud layer, decreasing the amount of radiation that reaches a planet’s surface. In the diffusive regime, radiation is more able to flow around clumps. Transport-regime conditions lie in between these two extremes.
@@
A straightforward approach to introduce inhomogeneites in reduced order techniques is to derive an effective opacity in the region with inhomogeneities. For example, the approach of Prinja & Olson (2005); Olson (2007) included an additional effective scattering opacity as well as a modification to the absorption opacity for an inhomogeneous medium for a purely absorptive medium (i.e. the scattering opacity is much lower than the absoprtion opacity). The effective absorption opacity ($\sigma_{\rm a,eff}$) is given by:
@@
where $\tilde{\sigma}=p_{1}\sigma_{a1}+p_{2}\sigma_{a2}$ is the average opacity weighted by the covering fraction, $\nu^{2}=p_{1}p_{2}(\sigma_{a1}-\sigma_{a2})^{2}$ is the unconditional variance in the opacity, and $\lambda_{c}=\lambda_{1}\lambda_{2}/(\lambda_{1}+\lambda_{2})$ is the geometric mean of the chord lengths, where the chord length is defined by the length of a chord covered by each material.
Using this prescription, we can calculate the effective absorption and scattering opacities as a function of the covering fraction of the clumps. Figure 1 shows the effective absorption and scattering opacities under this formalism as a function of the clump covering fraction. Although the effective absorption cross section does not increase until the clump covering fraction is large, the effective scattering term quickly approaches the value of the clump opacity. This scattering term can drastically lower the propagation speed of the radiation. It also provides a means to capture the angular redistribution from the clumps.
Figure 1: Effective absorption (solid) and scattering (dashed) opacities for three values of clump opacity (in ${\rm cm}^{2}\,{\rm g}^{-1}$) as a function of clump covering fraction. The opacity of the flow region is $10{\rm\,cm^{2}\,g^{-1}}$ and the density of the flow, clump regions are $0.05{\rm\,g\,cm^{-3}},5{\rm\,g\,cm^{-3}}$ respectively. The effective absorption opacity does not increase significantly until the covering fraction exceeds 0.8. The “OS” model corresponds to the  Owocki & Sundqvist (2018) formula (equation 3). However, the effective scattering opacity can be high if the clump opacity is high compared to the opacity in the flow region.
For stellar winds, Owocki & Sundqvist (2018) developed a mechanism that is a direct modification to the total opacity based on the porosity of the clumps. In this model, the effective opacity ($\sigma_{\rm eff}$) is given by:
$\frac{\sigma_{\rm eff}}{\sigma_{\rm clump}}=\frac{1-e^{-\tau_{h}}}{\tau_{h}}$ (3)
where $\tau_{h}=\langle\rho\sigma\rangle h$ where $h=(f_{\rm clump}-1)l_{\rm clump}$, $f_{\rm clump}=(l_{\rm sep}/l_{\rm clump})^{3}$, $l_{\rm clump}$ is the characteristic size of the clump and $l_{\rm sep}$ is the average separation of the clump. $f_{\rm clump}$ is equivalent to the covering factor of the clumps ($p_{2}$) used in equation 1. As the covering factor approaches 1, both models set the opacity to the clump opacity. In Owocki & Sundqvist (2018), the opacity of the flow region is assumed to be zero and, as the filling or covering factor approaches 0, the opacity converges to zero. To compare to the Olson (2007) solution, we set the final opacity to $\sigma_{\rm flow}+\sigma_{\rm eff}$ and include this opacity in Figure 1. This method focuses on capturing a mean opacity, but without an effective scattering term, is not able to capture the angular redistribution.
Neither of the above techniques include the shock heating and the effect it has on the radiation spectrum. More work is needed to compare the full physical effects of clumps on radiation flows.
Most of these past studies focused on transport-only solutions or solutions where radiation/hydrodynamic effects can be minimized. For example, Prinja & Olson (2005) tested their scheme using a radiative transfer code. Although the transfer could heat the material, these models were not coupled to a hydrodynamics package to model the subsequent expansion from the pressure gradient produced by the heating. As such, the solutions typically only consider the relative opacities of the different clumps. More detailed studies included the effects of the size scale of the clumps (with respect to the photon mean free path). This effect becomes increasingly important as we move from the diffusive to the free-streaming transport regimes.
3.2 Experiments and Scaling
A number of experiments studying radiation transport in inhomogeneous media already exist. These experiments are primarily focused on studying the propagation of radiation through this media. But a number of diagnostics have been improved to get at the effect of the inhomogeneities on the energy and angular distributions of the the propagating radiation. Dante X-ray spectrometers are often placed at the edge of the target to measure the coarse spectrum of the emerging radiation front (Sorce et al., 2006; Barnak et al., 2020). Absorption spectroscopy is being used to measure the atomic states, and hence temperature, of the material as the radiation front moves through the inhomogeneous medium Johns et al. (2023). These measurements allow experimentalists to more directly observe the response of the dense material to being heated by the radiation. These outflows drive shocks that can then be probed by the spectral measurements. Bear in mind, however, that these experiments can not be used to test opacity implementations in high-velocity media (e.g. expansion opacity recipes).
The scaling of experiments to applications has been discussed extensively in the literature (Drake, 2005; Falize et al., 2011; Keiter et al., 2017; Fryer et al., 2020b; Kurzer-Ogul et al., 2024). Radiation transport is well understood in extreme limits: in the optically-thick diffusion regime and the optically-thin free-streaming regime. What many radiation experiments probe is that intermediate transport regime. The uncertainties in numerical implementations of the Boltzmann transport equation are greatest in this regime. The primary experimental scaling parameter tying laboratory experiments to applications in this transport regime is the mean free path: e.g. number of mean free path across the target, number of mean free paths across a clump, number of mean free paths separating clumps. By using either light foam or gas materials in the flow region, experimentalists can vary these properties.
However, the goal is instead to study the reaction of the material to radiation. These experiments seek to test the hydrodynamics and transport schemes as well as the method used to couple these two physics models. The amount of outflow of a radiatively-heated dense clump/shell depends upon the characteristics of the clumps (e.g. densities, specific heats). But if we are more interested in testing numerical artifacts in numerics of radiation-hydrodynamics coupling: e.g. the outflow will depend on differences in the treatment between Lagrangian and Eulerian schemes (e.g. Kunishima & Onishi, 2018) and on effects of teleportation (or related numerical diffusion) in transport (Cheatham, 2010). With such studies direct scaling is not critical. Indeed, ideally the experiment is designed to exacerbate these effects so that we can more easily identify the needed improvements to the algorithms.
With these scaling issues in mind, we review a set of existing experiments. For code validation, it is important to understand the details of the experiments to both understand the uncertainties in the initial conditions, uncertainties in the physics assumptions and limitations of the diagnostics. The following detailed discussions give examples of the difficulties in producing accurate validation tests.
Shaoen et al. (2005) performed experiments investigating supersonic radiation propagation in a low-density heterogeneous medium. These experiments investigated radiation propagation in a low-density plastic foam target as well as in a Cu-doped foam. The Cu was added by means of small Cu particles and achieved roughly a 2.14% atomic fraction. The actual placement of the particles is difficult to control. The foams were viewed face-on with a trichromatic streaked x-ray spectrometer (TCS). The TCS consisted of a three-imaging-pinholes array and a three-transmission-grating array coupled with an x-ray streak camera. The TCS observed the emission from two different photon energies, 210 eV and 840 eV. A soft x-ray spectrometer observed the soft x-ray spectrum of the hohlraum, from which a temperature could be inferred. Finally, a transmission grating spectrometer was used to measure the time-integrated spectrum from either the hohlraum or the foam sample. A delay in the radiation breakout was observed when comparing the Cu-doped foam with the pure foam case for emission at 210 eV; however, this trend was reversed when comparing emission at 840 eV. The opacity for both the pure foam and the Cu-doped foam was estimated at each of these energies. This study concluded that differences in the opacities caused the different measured radiation flow speeds. But as we shall see here, the flow may also depend on the blow-off of the inclusions in the flow region.